OpenStudy (lcnettech):

Simplify the expression. 68 – 18 ÷ 3 • 6

OpenStudy (anonymous):

alright do this from left to right

OpenStudy (lcnettech):

ok

OpenStudy (anonymous):

68-18= 50/3=16.6666667x6

OpenStudy (lcnettech):

?

OpenStudy (anonymous):

=100

OpenStudy (lcnettech):

oh ok.

OpenStudy (anonymous):

do you understand or would you like me do explain it further?

OpenStudy (lcnettech):

was that the answer for this question?

OpenStudy (anonymous):

yes

OpenStudy (lcnettech):

well thanks

OpenStudy (lcnettech):

:)